OpenStudy (anonymous):

how you would use the elimination method to solve the following system of equations. Part 2: Provide the solution to the system. 8x + 5y = 3 7x – 3y = 10

OpenStudy (anonymous):

8x+5y=3-->(1) 7x-3y=10--->(2) ---------- (1)*7 56x+35y=21--->(3) (2)*8 56x-24y =80--->(4) -------------- (3)-(4) 56x+35y=21 -56x+24y=-80 --------------- 0 +59y=-59 =>y=-59/59 =>y=-1 put y=-1 in (1) 8x+5*(-1)=3 8x-5=3 8x=3+5 8x=8 =>x=8/8 =>x=1 Ans: x=1 and y=-1

OpenStudy (anonymous):

Thank You :)

OpenStudy (anonymous):

I hope that would help:)

OpenStudy (anonymous):

Ugh now i have to explain in complete sentence for those answers lol